Write an equation of the line containing the given point and parallel to given line. (5,-3); 6x-5y=7 The equation of the line is Y = ______.
My soulution is Y = 6/5x -33/5 The answer is Y = 6/5x -9 What makes my solution different from the answer..I have worked this several times. Any assistance is highly appreciated.

Maurice
email: mckinney7@earthlink.net

Answer by scott8148(6628)

not sure how you got -33/5

once you have the correct slope, you plug in the given point to find the y-intercept

y = (6/5)x + ? ___ (-3) = (6/5)(5) + ? ___ -9 = ?


looks like maybe you switched the x and y terms of the given point?